Circular No. 7351

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000E IN NGC 6951 G. Valentini, E. Di Carlo, D. Guidubaldi, and E. Brocato, Teramo

Observatory; and A. Di Paola, F. D’Alessio, R. Speziali, G. Li Causi, and F. Pedichini, Rome Observatory, report their discovery of a new supernova on CCD images obtained with the Teramo 0.72-m TNT telescope on Jan. 26.73 UT. SN 2000E is located at α = 20^{h}37^{m}13^{s}.8, δ = +66^{o}05′50′′.2 (equinox 2000.0), which is 6′′.3 west and 26′′.7 south of the center of NGC 6951, and 28′′ west and 24′′ south of SN 1999el (cf. IAUC 7288). Preliminary photometry of SN 2000E (near-infrared photometry obtained at the 1.1-m AZT-24 telescope at Campo Imperatore): Jan. 26.73, V = 14.30 ± 0.07; 26.78, J = 13.8 ± 0.1, H = 15.1 ± 0.1; 26.79, B = 15.78 ± 0.06. CCD frames taken on 1999 Dec. 18.78 show nothing at the location of 2000E, and the Digital Sky Survey does not show anything at that location. No object was detected in near-infrared images (J, H) obtained on Jan. 6.74.

M. Turatto, G. Galletta, and E. Cappellaro, Padova, report that preliminary inspection of a spectrogram (range 400–730 nm, resolution 1.4 nm) of SN 2000E, obtained with the 1.8-m telescope at Cima Ekar (+ Afosc) on Jan. 27.83 UT, shows a continuum with strong lines of Si II, S II, Si III, Fe III, and Ca II, indicating that SN 2000E is indeed a type-Ia event, a few days before maximum light. The expansion velocity deduced from the Si II lines is ∼ 11 000 km/s. Narrow absorption features of Na I D are present, indicating that the light from the supernova suffered consistent reddening both from the Milky Way and the parent galaxy.

SUPERNOVA 2000B IN NGC 2320 F. Colas, A. Fienga, and C. Buil, Pic du Midi Observatory, report that

a CCD spectrum (range 450–650 nm), obtained on Jan. 22.8 UT with the Pic du Midi 1-m telescope, shows that SN 2000B (cf. IAUC 7347) is of type Ia, near maximum brightness.

V723 CASSIOPEIAE D. Chochol and T. Pribulla, Skalnaté Pleso Observatory; O. Shemmer,

Wise Observatory, Tel Aviv University; A. Retter, Keele University; and S. Y. Shugarov, V. P. Goranskij, and N. A. Katysheva, Sternberg Astronomical Institute, Moscow, report: “Extensive BV R photometry of V723 Cas (N Cas 1995) during Dec. 1996 to Dec. 1999 reveals the nova’s binary period to be 0.69325(18) day. The full amplitude of the variations in the R band increased from 0.13 mag in 1998 to 0.35 mag in 1999.”

2000 January 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green